Find the equation of the line, given the point (−1, −5) and a slope of −8. Use point-slope form.
Answer by nerdybill(7384)

This site describes the "point-slope" form quite well:
http://www.purplemath.com/modules/strtlneq2.htm
.
Basically, the form is this:
y – y1 = m(x – x1)
.
From the problem, you were given:
a point (−1, −5) and a slope of −8
.
Plug the above into the "point-slope" form and then "rearrange":
y – y1 = m(x – x1)
y – (-5) = -8(x – (-1))
y + 5 = -8(x + 1)
y + 5 = -8x - 8
y = -8x - 13 (this is the answer)